OpenStudy (anonymous):

Solve each system graphically. Be sure to check your solution. If a system has an infinite number of solutions, use set−builder notation to write the solution set. If a system has no solution, state this.6x − 2y = 2, 9x − 3y = 1

OpenStudy (anonymous):

No solution. \[\frac{9}{6}(6x-2y)=2*\frac{9}{6}\to 9 x-3 y=3\]

OpenStudy (anonymous):

yes I know that but I am confused as to how to graph it for some reason I guess i for got how to graph my problem

OpenStudy (anonymous):

The lines are parallel. Plot attached. The blue line is the first equation, 6x − 2y = 2.